# The Value of AI Advice: Personalized and Value-Maximizing AI Advisors Are Necessary to Reliably Benefit Experts and Organizations

### v1

@@ 4.2 Producing TeamRules
Our framework can be used to derive different methods that generate ReV-AI advisors for different humans and contexts. We develop a rule-based prototype advisor, TeamRules (tr). Rules have been shown to offer an especially intuitive form of inherently interpretable advice for high-stakes decision-makers, given their transparent inner structures and expressiveness [31, 33, 34, 35, 36]. The purpose of the prototype is to first demonstrate the use of the framework to derive a method for generating ReV-AI advisors. Most importantly, the prototype allows us to study and characterize how the behaviors of different ReV-AI advisors’ produced by (tr) vary across contexts, and to study the ReV-AI advisors’ impact on value added to the human relative to that of alternatives.
@@
A rule $r$ is a logical expression made of conditional statements about a conjunction of a subset of feature values. $r$ covers an example $x_{i}$, denoted as $\mathbbm{C}(r,x_{i})=1$, if the logical expression evaluates to true given $x_{i}$.
@@
A rule set $R$ is a collection of rules. $R$ covers an example $x_{i}$, denoted $\mathbbm{C}(R,x_{i})=\mathbbm{1}\{\sum_{r\in R}\mathbbm{C}(r,x_{i})\geq 1\}$, if at least one rule $r$ in $R$ covers $x_{i}$.
Given a training set of instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, tr learns a set of rules $R=\{R^{+},R^{-}\}$ to advise the human, such that rules in $R^{+}$ or $R^{-}$ can be used to advise 1 or 0 on instances that satisfy them, respectively. Note that a distinct property of rule-based advice is that a given rule set $R$ can be produced so that it does not cover all instances. Following our framework, to minimize the expected loss from the human’s final decisions, a tr advisor selectively offers advice for an instance $x_{i}$ if $x_{i}$ is covered by a rule in $R$ (i.e., if $\mathbbm{C}(R,x_{i})=1$) and if the expected Total Team Loss (TTL) from offering the advice is lower than the expected TTL when advice is withheld. In summary, Figure 1b demonstrates TR’s advising process which consists first of providing possible advice determined by the rule set $R$ and by further assessing the expected added value of advice when an instance is covered by $R$.
@@
Selective Advising. Formally, tr’s selective advising at inference time is outlined in Equations (4) and (5). For any decision instance $x_{i}$ encountered during deployment, tr either offers advice or withholds advice which is equivalent to recommending the human’s independent decision $h_{i}$, given by: [^6] Note how Equation 1 from the ReV-AI advising framework is concretely operationalized to form tr’s rule-based advising process shown in Equation 4.
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbbm{C}(R^{+},x_{i})\wedge\psi(y,1,h_{i},\hat{p}(a),\alpha)\\ 0,&\text{if }\mathbbm{C}(R^{-},x_{i})\wedge\neg\mathbbm{C}(R^{+},x_{i})\wedge\psi(y,0,h_{i},\hat{p}(a),\alpha)\\ h_{i},&\text{otherwise}\end{cases}$ (4)
Where $\psi(y,y^{*},h,\hat{p}(a),\alpha)$ shown below is the indicator function that determines whether a prospective advice $y^{*}$ adds value in expectation, i.e., whether the expected loss from the advice ($\mathbbm{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$) is lower than the expected loss from withholding advice ($\mathbbm{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]$:
$\psi(y_{i},y^{*}_{i},h_{i},\hat{p}(a),\alpha)$ $=\mathbbm{1}\{\mathbbm{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$ (5)
$\text{\indent}<\mathbbm{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]\}$
To take expectation over the ground truth labels, $\mathbbm{E}_{y}$, we train a model $\hat{p}(y|x)$ using an XGBoost classifier [58] with hyperparameters selected using grid-search cross-validation [59]. In addition, the team loss from the human’s final decision for instance $x_{i}$, $\mathcal{L}(y_{i},\hat{y_{i}},h_{i},\hat{p}(a),\alpha)$, is shown in Equation (3). In our implementation of tr, we set the instance-level decision loss in Equation (3) to: $\mathcal{V}(y,d)=\lambda_{0}^{(1-y)}\lambda_{1}^{(y)}\mathbbm{1}(y\neq d)$, where $\lambda_{0}$ and $\lambda_{1}$ are the decision losses incurred from false positives and false negatives, respectively. All main results consider settings in which $\lambda_{0}=\lambda_{1}=1$, however, we also include additional results where $\lambda_{0}\neq\lambda_{1}$ in the Appendix.
@@
$c^{M}_{i}$ $=\underset{r\in R^{sign(\hat{y}_{i})}:\mathbbm{C}(r,x_{i})=1}{\max}\Bigg{[}\frac{\sum_{j=1}^{n}\big{(}\mathbbm{1}\{\mathbbm{C}(r,x_{j})\wedge\mathbbm{1}(y_{j}=\hat{y}_{i})\}\big{)}}{\sum_{k=1}^{n}\mathbbm{C}(r,x_{k})}\Bigg{]}$ (6)
@@
8:     for $i=1...n$ do
9:         Produce $\text{Produce }\hat{y}_{i}$, $c^{M}_{i}$triangleright According to Equations (4) and (6).
10:     ϵitalic-$\epsilon\leftarrow$ draw instance at index $i$ with probability: $\frac{\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\alpha\mathbbm{1}\{\mathcal{V}(y_{i},1-y_{i})\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})>\alpha\}}{\sum_{j=1}^{n}\bigg{(}\mathcal{L}(y_{j},\hat{y}_{j},h_{j},\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j}),\alpha)-\alpha\mathbbm{1}\{\mathcal{V}(y_{j},1-y_{j})\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j})>\alpha\}\bigg{)}}$
11:     if ($\mathbbm{C}(R_{t-1},x_{\epsilon})$ and ($\hat{y}_{\epsilon}\neq y_{\epsilon}$)) or (($h_{\epsilon}\neq\hat{y}_{\epsilon}$) and ($\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})<\alpha$)) then
@@
18:     for $i=1...n$ do
19:         Produce $\text{Produce }\hat{y}^{*}_{i}$, $c^{*M}_{i}$triangleright Using updated rule-set $R_{t}$
@@
22:     if $\exp\left({\frac{\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}^{*}_{i},h_{i},\hat{p}(a_{i}|c^{*M}_{i},c^{H}_{i},\hat{y}^{*}_{i}\neq h_{i}),\alpha)}{C_{0}^{\frac{t}{T}}}}\right)\leq$ random$([0,1])$ then
@@ 4.3 Optimization Approach
We develop a training algorithm based on the core idea of simulated annealing to learn a rule set $R$ that minimizes the team’s loss in Equation (3). Simulated annealing is well-suited for our discrete space, and has been used successfully in previous rule-based models [60, 61]. The complete algorithm is shown in Algorithm 1. The optimization procedure takes as input the set of training instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, along with the cost of reconciling a piece of contradictory advice $\alpha$ and the human discretion model $\hat{p}(a)$. Initially, a set $\Gamma$ of $\beta_{0}$ candidate rules is generated using an off-the-shelf rule mining algorithm, such as FP-growth [62]. We follow prior work and established practices to set the hyperparameters which we discuss in Appendix A.3. The initial rule set $R_{o}$ is set to an empty set so that decisions are made exclusively by the human. At each subsequent iteration $t$, a new rule set is proposed by either adding or dropping a rule to the current solution. This new set is accepted with a probability that depends on the quality of the proposed solution — i.e., the lower the total team loss (TTL) across all training instances, the higher the likelihood of acceptance. The probability also depends on the temperature, which decreases over iterations — i.e., the higher the temperature, the greater the likelihood of acceptance. As a result, the algorithm explores more aggressively at the beginning of the search, being more likely to accept solutions even if they are worse than the current one to thoroughly explore the space. Towards the end, it becomes more conservative, accepting only solutions that improve the existing rule set. This approach allows the algorithm to progressively minimize the TTL while avoiding getting stuck in local minima.

### v2

@@ 3.1 Producing TeamRules
Our framework can be used to derive different methods that generate ReV-AI advisors for different humans and contexts. We develop a rule-based prototype advisor, TeamRules (tr). Rules have been shown to offer an especially intuitive form of inherently interpretable advice for high-stakes decision-makers, given their transparent inner structures and expressiveness (rudin2019stop, ; wang2021scalable, ; rich, ; BALAGOPAL2021102101, ; edit_machine, ). The purpose of the prototype is to first demonstrate the use of the framework to derive a method for generating ReV-AI advisors. Most importantly, the prototype allows us to study and characterize how the behaviors of different ReV-AI advisors produced by (tr) vary across contexts, and to study the impact of ReV-AI advisors on value added to the human relative to that of alternatives.
@@
A rule $r$ is a logical expression made of conditional statements about a conjunction of a subset of feature values. $r$ covers an example $x_{i}$, denoted as $\mathbb{C}(r,x_{i})=1$, if the logical expression evaluates to true given $x_{i}$.
@@
A rule set $R$ is a collection of rules. $R$ covers an example $x_{i}$, denoted $\mathbb{C}(R,x_{i})=\mathbb{I}\{\sum_{r\in R}\mathbb{C}(r,x_{i})\geq 1\}$, if at least one rule $r$ in $R$ covers $x_{i}$.
Given a training set of instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, tr learns a set of rules $R=\{R^{+},R^{-}\}$ to advise the human, such that rules in $R^{+}$ or $R^{-}$ can be used to advise 1 or 0 on instances that satisfy them, respectively. Note that a distinct property of rule-based advice is that a given rule set $R$ can be produced so that it does not cover all instances. Following our framework, to minimize the expected loss from the human’s final decisions, a tr advisor selectively offers advice for an instance $x_{i}$ if $x_{i}$ is covered by a rule in $R$ (i.e., if $\mathbb{C}(R,x_{i})=1$) and if the expected Total Team Loss (TTL) from offering the advice is lower than the expected TTL when advice is withheld. In summary, Figure 3 demonstrates TR’s advising process which consists first of providing possible advice determined by the rule set $R$ and by further assessing the expected added value of advice when an instance is covered by $R$.
@@
Selective Advising. Formally, tr’s selective advising at inference time is outlined in Equations (4) and (3.1). For any decision instance represented by feature values $x_{i}$ encountered during deployment, tr either offers advice or withholds advice which is equivalent to recommending the human’s independent decision $h_{i}$, given by: [^4] Note how Equation 1 from the ReV-AI advising framework is concretely operationalized to form tr’s rule-based advising process shown in Equation 4.
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbb{C}(R^{+},x_{i})\wedge\psi(y,1,h_{i},\hat{p}(a),\alpha)\\ 0,&\text{if }\mathbb{C}(R^{-},x_{i})\wedge\neg\mathbb{C}(R^{+},x_{i})\wedge\psi(y,0,h_{i},\hat{p}(a),\alpha)\\ h_{i},&\text{otherwise}\end{cases}$ (4)
Where $\psi(y,y^{*},h,\hat{p}(a),\alpha)$ shown below is the indicator function that determines whether a prospective advice $y^{*}$ adds value in expectation, i.e., whether the expected loss from the advice ($\mathbb{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$) is lower than the expected loss from withholding advice ($\mathbb{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]$:
$\psi(y_{i},y^{*}_{i},$ $h_{i},\hat{p}(a),\alpha)=$
$\mathbb{I}\{\mathbb{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$ (5)
$<\mathbb{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]\}$
To take expectation over the ground truth labels, $\mathbb{E}_{y}$, we train a model $\hat{p}(y|x)$ using an XGBoost classifier (xgboost, ) with hyperparameters selected using grid-search cross-validation (scikit-learn, ). In addition, the total team loss from the human’s final decision for decision instance at index $i$, $\mathcal{L}(y_{i},\hat{y_{i}},h_{i},\hat{p}(a),\alpha)$, is shown in Equation (3). In our implementation of tr, we set the instance-level decision loss in Equation (3) to: $\mathcal{V}(y,d)=\lambda_{0}^{(1-y)}\lambda_{1}^{(y)}\mathbb{I}(y\neq d)$, where $\lambda_{0}$ and $\lambda_{1}$ are the decision losses incurred from false positives and false negatives, respectively.
@@
$c^{M}_{i}=$ (6)
$\underset{r\in R^{sgn(\hat{y}_{i})}:\mathbb{C}(r,x_{i})=1}{\max}\Bigg[\frac{\sum_{j=1}^{n}\big(\mathbb{I}\{\mathbb{C}(r,x_{j})\wedge\mathbb{I}(y_{j}=\hat{y}_{i})\}\big)}{\sum_{k=1}^{n}\mathbb{C}(r,x_{k})}\Bigg]$
@@
8:  for i=1ni=1...n do
9:   $\text{Produce }\hat{y}_{i}$, $c^{M}_{i}$ $\triangleright$ According to Equations (4) and (6).
10:  $\epsilon\leftarrow$ draw instance at index $i$ with probability: $\frac{\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\alpha\mathbb{I}\{\mathcal{V}(y_{i},1-y_{i})\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})>\alpha\}}{\sum_{j=1}^{n}\bigg(\mathcal{L}(y_{j},\hat{y}_{j},h_{j},\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j}),\alpha)-\alpha\mathbb{I}\{\mathcal{V}(y_{j},1-y_{j})\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j})>\alpha\}\bigg)}$
11:  if ($\mathbb{C}(R_{t-1},x_{\epsilon})$ and ($\hat{y}_{\epsilon}\neq y_{\epsilon}$)) or $($$h_{\epsilon}\neq\hat{y}_{\epsilon}$) and ($\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})<\alpha$)) then
@@
18:  for i=1ni=1...n do
19:   $\text{Produce }\hat{y}^{*}_{i}$, $c^{*M}_{i}$ $\triangleright$ Using updated rule-set $R_{t}$
@@
22:  if $\exp\left({\frac{\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}^{*}_{i},h_{i},\hat{p}(a_{i}|c^{*M}_{i},c^{H}_{i},\hat{y}^{*}_{i}\neq h_{i}),\alpha)}{C_{0}^{\frac{t}{T}}}}\right)\leq$ random([0,1])([0,1]) then
@@ 3.2 Optimization Approach
We develop a training algorithm based on the core idea of simulated annealing to learn a rule set $R$ that minimizes the team’s loss in Equation (3). Simulated annealing is well-suited for our discrete space, and has been used successfully in previous rule-based models (sim_anneal, ; wang2017bayesian, ). The complete algorithm is shown in Algorithm 1. The optimization procedure takes as input the set of training instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, along with the cost of reconciling a piece of contradictory advice $\alpha$ and the human discretion model $\hat{p}(a)$. Initially, a set $\Gamma$ of $\beta_{0}$ candidate rules is generated using an off-the-shelf rule mining algorithm, such as FP-growth (wang_gaining_nodate, ). We follow prior work and established practices to set the hyperparameters which we discuss in Appendix A.3. The initial rule set $R_{o}$ is set to an empty set so that decisions are made exclusively by the human. At each subsequent iteration $t$, a new rule set is proposed by either adding or dropping a rule to the current solution. This new set is accepted with a probability that depends on the quality of the proposed solution — i.e., the lower the total team loss (TTL) across all training instances, the higher the likelihood of acceptance. The probability also depends on the temperature, which decreases over iterations — i.e., the higher the temperature, the greater the likelihood of acceptance. As a result, the algorithm explores more aggressively at the beginning of the search, being more likely to accept solutions even if they are worse than the current one to thoroughly explore the space. Towards the end, it becomes more conservative, accepting only solutions that improve the existing rule set. This approach allows the algorithm to progressively minimize the TTL while avoiding getting stuck in local minima.